What is the probability of this independent event. Drawing a red marble from a bag of 6 red and 4 blue marbles, replacing it, and then drawing a blue marble.
Answer by blimblim(1)

Drawing a red marble the chance is 6/10. After replacement of the red marble the chance of drawing a blue is 4/10. Since these are independent events you derive the result of both drawing the red marble and then drawing a blue marble by multiplying the two probabilies. Hence, the answer is
6/10 x 4/10 = 24/100 = 6/25 = 24%.
Hope this helps
Sincerely,
BlimBlim (the wizard)